A certain compound has a solubility of 169 g/L in a solvent.If the mixture contains 50 g of solute in a 300 cubic cm of solution, is the solution unsaturated? If yes, how many grams of solute should be added to make a saturated solution?
1
Expert's answer
2020-10-12T13:30:26-0400

169 / 1000 = 0.169

0.169 * 300 = 50.7 > 50; solution unsaturated

50.7 - 50 = 0.7; should be add 0.7 g;